OpenStudy (anonymous):

The function f(x) = 396(3)x represents the growth of a dragonfly population every year in a remote swamp. Pierre wants to manipulate the formula to an equivalent form that calculates every month, not every year. Which function is correct for Pierre's purposes? I don't get it, please help?

OpenStudy (boldjon):

ok gimme a sec

OpenStudy (mathmale):

This problem just might make more sense if you'd express the exponentiation properly. Correct: f(x) = 396(3)^x Use " ^ " to indicate exponentiation. Incorrect: f(x) = 396(3)x

OpenStudy (anonymous):

Sorry ;(

OpenStudy (mathmale):

I'll demonstrate a parallel (comparable) situation: \[A=P(1+r)^t\]

OpenStudy (anonymous):

Since its by month would I divide the whole equation by 12?

OpenStudy (mathmale):

is the formula used to calculate the Amount of $ that you will have after t years. If you're interested in monthly compounding, then the appropriate formula is\[A=P(1+\frac{ r }{ n })^{nt}\]

OpenStudy (mathmale):

where n is the number of times per year that interest is paid (compounded). So yes, you're on the right track.

OpenStudy (anonymous):

would we use the second formula then?

OpenStudy (mathmale):

As explained, the 2nd formula was meant to be an example. It's a modification of the first formula for Amount. Can you modify f(x) = 396(3)^x to reflect monthly (rather than annual) increases?

OpenStudy (anonymous):

i figured it out

OpenStudy (anonymous):

want to help me with some others?